Week 1: Dosage Calculations Practice Questions 1) Calculate the fluid intake in milliliters (mL) for the following food items. 3 oz grape juice 6 oz broth 9 oz coffee 12 oz soft drink TOTAL = _____ mL a. 243 b. 1,800 c. 420 d. 900 2) How much IV fluid would you document as fluid intake if 350 mL is left in a 1,000-mL IV bag? _______________ a. 550 mL b. 650 mL c. 350 mL d. 750 mL 3) How much IV fluid would you document as fluid intake if 625 mL is left in a 1,000-mL IV bag? _______________ a. 375 mL b. 275 mL c. 400 mL d. 625 mL 4) Order: Lanoxin 0.25 mg Available: 5) Order: Hydrochlorothiazide 0.05 g Available: 6) Order: Thorazine 75 mg Available: 7) Order: Lanoxin 125 mcg Available: 8) Order: Mycostatin 300,000 units Available: 9) Order: Epivir 150 mg Available: 10) Order: Prozac 50 mg Available:
Answer #1

Week 1:

Dosage Calculations Practice Questions 1) Calculate the fluid intake in milliliters (mL) for the following food items. 3 oz grape juice 6 oz broth 9 oz coffee 12 oz soft drink TOTAL = d. 900mL

Step 1:

1 oz = 30mL

Therefore,

3 oz grape juice = 3 x 30 = 90mL

6 oz broth = 6 x 30 = 180mL

9 oz cofee = 9 x 30 = 270mL

12 oz soft drink = 12 x 30 = 360mL.

Step 2:

Total mL = 90 + 180 + 270 +360 = 900mL.

2) How much IV fluid would you document as fluid intake if 350 mL is left in a 1,000-mL IV bag? b. 650mL

Total fluid = 1000mL

Remaining fluid = 350mL

Total intake = 1000 - 350 = 650mL.

3) How much IV fluid would you document as fluid intake if 625 mL is left in a 1,000-mL IV bag? a. 375mL

Total fluid = 1000mL

Remaining fluid left in the bag = 625mL

Total intake = 1000 - 625 = 375mL.

4) Order: Lanoxin 0.25 mg

Available: Digoxin 0.25mg,

5) Order: Hydrochlorothiazide 0.05 g

Hydrochlorothiazide 50mg

Available: 12.5mg, 25mg, 50mg.

6) Order: Thorazine 75 mg Available:

Available: Chlorpromazine 30mg, 75mg, 150mg, 25mg, 100mg.

7) Order: Lanoxin 125 mcg

Available: Digoxin 0.125mg

8) Order: Mycostatin 300,000 units

Available: Nystatin 100,000units per gram

9) Order: Epivir 150 mg

Available: Epivir 150mg

10) Order: Prozac 50 mg

Available: Prozac10mg, 20mg, 40mg.